# Physical-regularized Hierarchical Generative Model for Metallic Glass Structural Generation and Energy Prediction

### v1

@@ Abstract
[…]
1 Introduction
@@
Atomistic simulations of disordered materials—glasses in particular—are computationally intensive. The absence of long-range periodicity in these systems prevents the use of compact unit-cell descriptions, making new configurations have to be generated from first principles.[1, 2] While generative AI has been successfully applied to crystalline solids [3, 4, 5, 6] and molecules by using structure representations like the Crystallographic Information File (CIF) [7, 8] and Simplified Molecular Input Line Entry System (SMILES)[9, 10, 11, 12], yet, no equivalent representation exists for inorganic glasses. Consequently, researchers rely on local geometric fingerprints—such as smooth overlap of atomic positions (SOAP) kernels [13, 14, 15], or atom-centered symmetry functions (ACSF) [16] —which capture only short-range order and omit system-level geometry and related properties. Generative models trained solely on these local features often propose structures that appear locally plausible yet not satisfying global property constraints, hindering the discovery of new structures and materials.
@@
In addition, glasses are inherently non-equilibrium systems with complex, high-dimensional configurational energy landscapes[17]. Exploring these landscapes to identify realistic glassy structures is computationally intensive, as such structures occupy only a tiny subset of all possible configurations. Accurate generation of glass structures therefore requires the modeling to be aware of their configurational energy. Traditionally, molecular dynamics (MD) simulations have been the primary tool, where atomic interactions are explicitly simulated. By replicating the melt-quench process using MD, specific regions of the energy landscape are explored and yields structures resembling experimentally produced glasses. However, due to computational constraints, MD quenching rates are typically 6–9 orders of magnitude faster than the experimental rates[18, 19]. This results in structures with lower thermodynamic stability and altered properties[20]. This timescale mismatch is also a major challenges for simulating key glass behaviors such as relaxation and aging[21, 22].
@@
To address these challenges, graph-based representations combined with generative frameworks capable of predicting configurational energy offer a promising approach for modeling disordered systems like glasses. In these representations, atoms are treated as nodes and interatomic distances serve as invariant edge features. Through symmetric aggregation, such models inherently respect E(3) symmetries—translation, rotation, and permutation invariance—enabling them to encode essential physical constraints directly into the architecture.[23, 24] This built-in invariance allows for the compression of high-dimensional atomic configurations into a compact, continuous latent space, significantly reducing computational overhead while retaining critical structural information. An explicit latent space is particularly useful for studying glass systems. Traversing this latent space enables the generation of new glassy structures and systematic exploration of their thermodynamic properties.
@@
Despite the strong predictive performance of graph neural networks (GNNs) for atomic-scale property estimation [25], embedding them within a generative modeling framework remains nontrivial. Hierarchical latent-space variational autoencoders (H-VAEs) [26, 27, 28] overcome this by learning multiscale latent representations: high-level variables capture global composition and medium to long-range features, while low-level variables encode local geometry. This architecture enhances expressiveness, mitigates posterior collapse, and produces physically plausible atomistic structures.
@@
Here, we introduce GlassVAE, a novel hierarchical variational autoencoder specifically designed to compress high-dimensional atomic graphs of disordered systems into a structured, low-dimensional latent space. GlassVAE’s effectiveness stems from two key innovations: (i) Physics-Informed Regularization: A radial-distribution-function (RDF) loss ensures the preservation of essential short- and medium-range structural order [29, 30], while an energy-regression loss aligns the latent embeddings with thermodynamic reality. (ii) Specialized Hierarchical Latent Structure: A graph-level latent variable captures global characteristics such as composition and the overall energy landscape, while a distinct edge-level latent variable refines local geometric details. This strategic division of responsibilities within the hierarchy effectively balances global energetic considerations with local structural fidelity, without unduly inflating the dimensionality of the latent space.
@@
In this work, we demonstrate a prototype GlassVAE on the CuZr metallic glass system. This system is selected to illustrate both the potential and the complexity of disordered alloys.[31] These amorphous metals combine the high strength of crystalline metals with the formability of polymers. [32] Their non-crystalline atomic arrangement underlies exceptional mechanical, thermal, and corrosion-resistance properties, driving their application in diverse fields including aerospace components, biomedical implants, and advanced sports equipment.[33, 34] Despite extensive study, the details of their structure-property relationship are still not well understood. [35, 36, 37] Through theoretical analysis and experiments on CuZr metallic glass, we demonstrate that GlassVAE accurately reproduces potential-energy distributions, achieves low root-mean-square error (RMSE) in reconstructing atomic positions, and, crucially, generates novel configurations that simultaneously adhere to RDF profiles and energetic constraints. By effectively unifying geometric and energetic priors within a single, tractable latent space, GlassVAE establishes the first end-to-end generative framework specifically tailored for disordered inorganic systems. This work opens a new avenue for AI-guided discovery and design of complex glassy materials.
2 Related Work
@@
Generative AI in Materials Science  GenAI models—including Generative Adversarial Networks (GANs) [38, 39], Variational Autoencoders (VAEs)[40, 41], autoregressive models like Transformers [42], and diffusion models [23, 24]—have become powerful tools in materials design. These approaches are increasingly applied to sample novel compositions, predict structures, and estimate properties, often within a unified computational framework [43, 44]. Significant breakthroughs have been demonstrated, particularly for crystalline materials, [45, 8, 46] and for organic molecules, enabling de novo drug design and property optimization[47, 40, 48]. Despite these successes, most generative efforts remain confined to systems exhibiting well-defined order, such as the periodic arrangements in crystals or the discrete bonding patterns in molecules. While generative techniques have been applied to glasses, [49, 50]these studies have largely focused on predicting macroscopic properties, generating microstructure morphologies, rather than atomic-level configurations [2].
@@
Graph Neural Networks  Graph neural networks (GNNs) have become indispensable tools in molecular and materials simulation, enabling accurate prediction of energies, forces, and other properties [51, 23, 52]. Among all, Wang et al., combined GNNs with swap Monte Carlo to inversely design Cu–Zr metallic glasses with enhanced plastic resistance [53], and Li et al., applied GNNs to predict energy barriers and interpret the behavior of metallic glasses [25]. Yoshikawa et al., integrated self-attention into a GNN for structural classification of glass-forming liquids [54], while Bapst et al., trained a GNN solely on initial particle positions to predict long-time glass dynamics [55]. Although these studies demonstrate the power of GNNs for property prediction and classification in glassy systems, none simultaneously address structural generation. Our approach fills this gap by learning an internal generative mechanism that bridges state-of-the-art molecular-graph generative models with physically consistent atomic reconstructions.
@@
Physical-informed Descriptors  Global, physics-informed descriptors—such as Crystallographic Information Files (CIFs) for crystals [7, 8] and SMILES for organic molecules [11, 9, 10, 56]—require long-range order and thus fail for glasses. Instead, local geometric fingerprints like smooth overlap of atomic positions (SOAP) kernels [15, 13, 14] or atom-centered symmetry functions (ACSF) [16] capture only short-range structure and omit system-level properties such as total potential energy. The radial distribution function (RDF) encodes short- and medium-range order in disordered materials. For example, Krykunov et.al., introduced atomic property–weighted RDFs with analytic Gaussian kernels that both outperform and accelerate Bag-of-Bonds for atomization energy prediction [57], and Watanabe et.al., developed an on-the-fly RDF-based sampling method to flag anomalous configurations and build robust water potentials [58]. To our knowledge, we are among the first to integrate an RDF-based loss into a variational graph model for glasses, combining RDF-driven regularization with graph representations to capture both local geometry and global energetics.
3 Physical-regularized Hierarchical Graph Variational Autoencoder
@@
In this work, we introduce GlassVAE, a generative framework for disordered metallic–glass configurations that jointly captures local atomic geometry and global energetics. Our approach addresses three core tasks: (i) Reconstruction: Recover atom types, interatomic displacements and distances, and absolute positions; (ii) Energy prediction: Estimate the total potential energy of each configuration; (iii) Generation: Sample new, physically realistic atomic structures from the learned latent distribution. Formally, each molecular-dynamics snapshot is represented by $(\mathbf{R},\mathbf{t},E)$, where $\mathbf{R}=\{\mathbf{r}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}\quad\text{(positions in a periodic box }\mathbf{L}=[L_{x},L_{y},L_{z}]\text{),}$, $\mathbf{t}$ is the one-hot encoding of atom types, and $E$ is the total potential energy. We learn $\mathcal{E}:(\mathbf{R},\mathbf{t},E)\;\mapsto\;\bm{z},\quad\mathcal{D}:\bm{z}\;\mapsto\;(\mathbf{R},\mathbf{t}),\quad\text{and}\quad\mathcal{P}:\bm{z}\;\mapsto\;\mathbf{R}$ so that the latent variable $\bm{z}$ retains all essential physical information for faithful reconstruction, accurate energy estimation, and the generation of novel, physically plausible structures.
3.1 Graph Representation and Invariance
@@
Physical interactions in materials depend only on atomic species and their relative arrangements, not on ordering, absolute position, or orientation. Therefore, atomistic representations and predictions must be invariant under permutation, translation, and rotation—swapping identical atoms, shifting all positions by $\Delta t$, or applying any orthogonal rotation $R$ leaves interatomic distances (and thus all physical properties) unchanged. To enforce these invariance, we convert each point cloud into an attributed graph $\mathcal{G}=(V,E)$, where Nodes $i\in V$ carry one-hot features $t_{i}$ encoding atomic species i.e., $x_{i}\in\{0,1\}$; and Edges $(i,j)\in E$ connect atoms within a cutoff distance (under periodic boundaries) and carry attributes
@@
A graph neural network with symmetric (e.g., mean) message-passing layers aggregates these features, guaranteeing invariance to node ordering and rigid-body transformations. We further support these invariance claims with a formal theorem.
Theorem 1 (Invariance of the Graph Representation).
Let an atomic configuration be given by positions =$\mathcal{R}=\{\mathbf{r}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{3}$ and one-hot node features $\{\mathbf{t}_{i}\}$. Construct a graph $G=(V,E)$ by connecting all pairs with $\|\mathbf{r}_{i}-\mathbf{r}_{j}\|_{2}\leq\mathrm{CUTOFF}$, and assign each edge the attribute
$\mathbf{a}_{ij}=\bigl{(}\Delta\bm{r_{ij}},\|\mathbf{r}_{i}-\mathbf{r}_{j}\|_{2}\bigr{)},\quad\text{where}\ \Delta\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}.$
Suppose the graph neural network updates and pools features using a symmetric (permutation-invariant) aggregation function $S$ (e.g. mean or sum). Then the resulting representation is invariant to
@@
so all edge-wise inputs to the network are identical.
@@
so all distance-based edge features are preserved.
@@
3.2 Model Architecture
To jointly capture the intricate local geometry and global properties (e.g., energies) of disordered metallic glass configurations, we propose a hierarchical graph variational autoencoder (GlassVAE). Figure 1 sketches its architecture. Sec. 3.1 states how we build a graph starting from an atomic configuration (also see Sec. A.2, A.3 for more details). From graph representations, we have
Dual-path Encoder  The dual-path encoder that distills two complementary summaries. The graph-level encoding applies $L$ layers of message passing to embed local neighbourhoods yielding a graph vector that two linear heads convert into the variational parameters $\mu$ and $\log\sigma^{2}$. We could then reparameterize them into a decoder head $z\sim\mu+\sigma$. In parallel, the edge path feeds all $a_{ij}$ through a residual stack of $L_{e}$ blocks and averages the refined edge features into an edge descriptor $\bm{s}$.
@@ 3.2 Model Architecture
Property Predictor  A separate fusion head consumes the concatenated vector $\Phi(\bm{z}\!\parallel\!\bm{s})$ to predict the total energy $\hat{E}$, allowing the model to learn structure–property relationships directly from the latent space.
@@
$\mathcal{L}=\;\alpha_{\text{node}}\,\mathcal{L}_{\text{node}}+\alpha_{\text{edge}}\,\mathcal{L}_{\text{edge}}+\alpha_{\text{energy}}\,\mathcal{L}_{\text{energy}}+\beta_{\text{KL}}\,D_{\text{KL}}+\alpha_{\text{RDF}}\,\mathcal{L}_{\text{RDF}}.$ (1)
@@
$\mathcal{L}_{\mathrm{node}}$ is cross-entropy between ground-truth one-hot atom types and predicted class probabilities; $\mathcal{L}_{\mathrm{edge}}$ is MSE on predicted distances plus a cosine-similarity penalty to preserve directional consistency of edge features; $\mathcal{L}_{\mathrm{energy}}$ is MSE between predicted and true potential energies; $D_{\mathrm{KL}}$ is Kullback-Leibler (KL) divergence; $\mathcal{L}_{\mathrm{RDF}}$ is $\ell_{2}$ distance between soft histograms of predicted and true pairwise distance distributions.Details about how we calculated the loss can be refered to Sec.A.3.
3.3 Generation
Generation is one of the key features of GlassVAE: its structured latent space enables both random and energy-conditioned synthesis of new, physically realistic configurations through the decoder. Here, we outline the two modes of operation. We also list the results of generation in Sec. 4.3.
@@ 3.3 Generation
$\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\text{and set}\ \mathbf{z}=\bm{\mu}(x)+\bm{\sigma}(x)\,\odot\,\gamma\bm{\epsilon},$
where $\odot$” denotes element-wise multiplication. This turns sampling into a differentiable operation with respect to $\bm{\mu}$ and $\bm{\sigma}$. here $\bm{\epsilon}$ is a small perturbation. Then the sampled new structures could be decoded through $\hat{x}=D_{1}(z)\ \text{and}\ \hat{a}=D_{2}(z)$.
Conditional generation. To constrain the generated energy to $[E_{\min},E_{\max}]$, we could use conditional generation. Formally, we assume a standard normal prior $p(z)\sim\mathcal{N}(0,I)$. To enforce $\hat{E}(z)\in[E_{\min},E_{\max}]$, we initialize $z$ at the encoder’s posterior mean and minimize
$\mathcal{L}(z)=\bigl{[}\max(E_{\min}-\hat{E}(z),0)\bigr{]}^{2}+\bigl{[}\max(\hat{E}(z)-E_{\max},0)\bigr{]}^{2}+\lambda\|z\|_{2}^{2}$
@@
3.4 Theoretical Contribution of Physics-Informed Regularizer
Physics-informed loss terms (i.e. energy and RDF in our case, Eq.1) do more than improve reconstruction accuracy—they can associate the latent variables to physically admissible manifolds. Below we formalize a theoretical explanation: imposing an operator-based regularizer $\|\mathcal{A}\bm{u}-v\|_{L^{2}}$ induces a bound on the latent test error that decays with the training error and the quadrature resolution. In other words, every unit of training progress achieved by a physics-informed term translates directly into a proportionate tightening of the latent-space generalization gap (see full proof in Sec. A.4).
Functional setting. Let $\mathcal{U}=W^{s,q}(\mathbb{R}^{3})$ and $\mathcal{V}=L^{p}(\mathbb{R}^{3})$ with $1\!\leq\!p,q\!\leq\!\infty$ and $s\!\geq\!0$. Given a neural approximation $\bm{u}(\cdot;\theta)\!\in\!\mathcal{U}$, the operator $\mathcal{A}:\mathcal{U}\!\to\!\mathcal{V}$ extracts a physical observable (e.g. predicted RDF, energy density), while $v\!\in\!\mathcal{V}$ encodes the target dictated by first-principles or simulation data.
Assumption 1 (Operator stability).
There exists $C_{1}>0$ such that $\|\bm{u}-\bm{u}^{\prime}\|_{\mathcal{U}}\leq C_{1}\|\mathcal{A}\bm{u}-\mathcal{A}\bm{u}^{\prime}\|_{\mathcal{V}}$ for all $\bm{u},\bm{u}^{\prime}\in\mathcal{U}$.
Assumption 2 (Quadrature accuracy).

### v2

# Physical‑regularized Hierarchical Generative Model for Metallic Glass Structural Generation and Energy Prediction

@@ Abstract
Disordered materials such as glasses, unlike crystals, lack long‑range atomic order and have no periodic unit cells, yielding a high‑dimensional configuration space with widely varying properties. The complexity not only increases computational costs for atomistic simulations but also makes it difficult for generative AI models to deliver accurate property predictions and realistic structure generation. In this work, we introduce GlassVAE, a hierarchical graph variational autoencoder that uses graph representations to learn compact, translation‑, and permutation‑invariant embeddings of atomic configurations. The resulting structured latent space not only enables efficient generation of novel, physically plausible structures but also supports exploration of the glass energy landscape. To enforce structural realism and physical fidelity, we augment GlassVAE with two physics‑informed regularizers: a radial distribution function (RDF) loss that captures characteristic short‑ and medium‑range ordering and an energy regression loss that reflects the broad configurational energetics. Both theoretical analysis and experimental results highlight the critical impact of these regularizers. By encoding high‑dimensional atomistic data into a compact latent vector and decoding it into structures with accurate energy predictions, GlassVAE provides a fast, physics‑aware path for modeling and designing disordered materials.
Code & Datahttps://github.com/EricCH97/GlassVAE
Introduction
@@
Atomistic simulations of disordered materials—glasses in particular—are computationally intensive. The absence of long-range periodicity in these systems prevents the use of compact unit-cell descriptions, making new configurations have to be generated from first principles.(bamerMolecularMechanicsDisordered2023; jungRoadmapMachineLearning2025) While generative AI has been successfully applied to crystalline solids (court3DInorganicCrystal2020; liGenerativeDesignCrystal2025; longConstrainedCrystalsDeep2021; nouiraCrystalGANLearningDiscover2019) and molecules by using structure representations like the Crystallographic Information File (CIF) (flam-shepherdLanguageModelsCan2023; zeniMatterGenGenerativeModel2024) and Simplified Molecular Input Line Entry System (SMILES)(oboyle_deepsmiles_2018; skinnider_invalid_2024; weininger_smiles_1988; sun_ice_2024), yet, no equivalent representation exists for inorganic glasses. Consequently, researchers rely on local geometric fingerprints—such as smooth overlap of atomic positions (SOAP) kernels (de_comparing_2016; lin_expanding_2024; bartok_representing_2013), or atom-centered symmetry functions (ACSF) (mudassir_systematic_2022) —which capture only short‑range order and omit system‑level geometry and related properties. Generative models trained solely on these local features often propose structures that appear locally plausible yet not satisfying global property constraints, hindering the discovery of new structures and materials.
@@
In addition, glasses are inherently non-equilibrium systems with complex, high-dimensional configurational energy landscapes(debenedetti_supercooled_2001). Exploring these landscapes to identify realistic glassy structures is computationally intensive, as such structures occupy only a tiny subset of all possible configurations. Accurate generation of glass structures therefore requires the modeling to be aware of their configurational energy. Traditionally, molecular dynamics (MD) simulations have been the primary tool, where atomic interactions are explicitly simulated. By replicating the melt-quench process using MD, specific regions of the energy landscape are explored and yields structures resembling experimentally produced glasses. However, due to computational constraints, MD quenching rates are typically 6–9 orders of magnitude faster than the experimental rates(binder_molecular_2004). This results in structures with lower thermodynamic stability and altered properties(yu_structural_2021). This timescale mismatch is also a major challenges for simulating key glass behaviors such as relaxation and aging(yu_stretched_2015; berthier_modern_2023).
@@
To address these challenges, graph-based representations combined with generative frameworks capable of predicting configurational energy offer a promising approach for modeling disordered systems like glasses. In these representations, atoms are treated as nodes and interatomic distances serve as invariant edge features. Through symmetric aggregation, such models inherently respect E(3) symmetries—translation, rotation, and permutation invariance—enabling them to encode essential physical constraints directly into the architecture.(satorras_en_2022; bao_equivariant_2023) This built-in invariance allows for the compression of high-dimensional atomic configurations into a compact, continuous latent space, significantly reducing computational overhead while retaining critical structural information. An explicit latent space is particularly useful for studying glass systems. Traversing this latent space enables the generation of new glassy structures and systematic exploration of their thermodynamic properties.
@@
Despite the strong predictive performance of graph neural networks (GNNs) for atomic-scale property estimation (liPredictingInterpretingEnergy2024), embedding them within a generative modeling framework remains nontrivial. Hierarchical latent-space variational autoencoders (H-VAEs) (klushynLearningHierarchicalPriors2019; vahdatNVAEDeepHierarchical2020; yingHierarchicalGraphRepresentation2019) overcome this by learning multiscale latent representations: high‑level variables capture global composition and medium to long-range features, while low‑level variables encode local geometry. This architecture enhances expressiveness, mitigates posterior collapse, and produces physically plausible atomistic structures.
@@
Here, we introduce GlassVAE, a novel hierarchical variational autoencoder specifically designed to compress high-dimensional atomic graphs of disordered systems into a structured, low-dimensional latent space. GlassVAE’s effectiveness stems from two key innovations: (i) Physics-Informed Regularization: A radial-distribution-function (RDF) loss ensures the preservation of essential short- and medium-range structural order (rapaport_art_2004; allen2017computer), while an energy-regression loss aligns the latent embeddings with thermodynamic reality. (ii) Specialized Hierarchical Latent Structure: A graph-level latent variable captures global characteristics such as composition and the overall energy landscape, while a distinct edge-level latent variable refines local geometric details. This strategic division of responsibilities within the hierarchy effectively balances global energetic considerations with local structural fidelity, without unduly inflating the dimensionality of the latent space.
@@
In this work, we demonstrate a prototype GlassVAE on the CuZr metallic glass system. This system is selected to illustrate both the potential and the complexity of disordered alloys.(pauly_transformation-mediated_2010) These amorphous metals combine the high strength of crystalline metals with the formability of polymers. (schroers_thermoplastic_2011) Their non‑crystalline atomic arrangement underlies exceptional mechanical, thermal, and corrosion‑resistance properties, driving their application in diverse fields including aerospace components, biomedical implants, and advanced sports equipment.(trexler_mechanical_2010; bansal_chapter_1986) Despite extensive study, the details of their structure-property relationship are still not well understood. (starr_what_2002; cao_structural_2018) Through theoretical analysis and experiments on CuZr metallic glass, we demonstrate that GlassVAE accurately reproduces potential-energy distributions, achieves low root-mean-square error (RMSE) in reconstructing atomic positions, and, crucially, generates novel configurations that simultaneously adhere to RDF profiles and energetic constraints. By effectively unifying geometric and energetic priors within a single, tractable latent space, GlassVAE establishes the first end-to-end generative framework specifically tailored for disordered inorganic systems. This work opens a new avenue for AI-guided discovery and design of complex glassy materials.
Related Work
@@
Generative AI in Materials Science GenAI models—including Generative Adversarial Networks (GANs) (nouira_crystalgan_2019; xuGenerativeAdversarialNetworks2023), Variational Autoencoders (VAEs)(gomez-bombarelli_automatic_2018; luo_deep_2024), autoregressive models like Transformers (wang_compositionally_2021), and diffusion models (satorras_en_2022; bao_equivariant_2023)—have become powerful tools in materials design. These approaches are increasingly applied to sample novel compositions, predict structures, and estimate properties, often within a unified computational framework (schleder_dft_2019). Significant breakthroughs have been demonstrated, particularly for crystalline materials, (xie_graph_2019; zeniMatterGenGenerativeModel2024; ren_invertible_2022) and for organic molecules, enabling de novo drug design and property optimization(gomez-bombarelli_automatic_2018; hoogeboom_equivariant_2022). Despite these successes, most generative efforts remain confined to systems exhibiting well-defined order, such as the periodic arrangements in crystals or the discrete bonding patterns in molecules. While generative techniques have been applied to glasses, (li_inverse_2024; cassar_designing_2021)these studies have largely focused on predicting macroscopic properties, generating microstructure morphologies, rather than atomic-level configurations (jungRoadmapMachineLearning2025).
@@
Graph Neural Networks Graph neural networks (GNNs) have become indispensable tools in molecular and materials simulation, enabling accurate prediction of energies, forces, and other properties (gilmer_neural_2017; satorras_en_2022; sanchez-gonzalez_learning_2020). Among all, Wang et al., combined GNNs with swap Monte Carlo to inversely design Cu–Zr metallic glasses with enhanced plastic resistance (wang_inverse_2021), and Li et al., applied GNNs to predict energy barriers and interpret the behavior of metallic glasses (liPredictingInterpretingEnergy2024). Yoshikawa et al., integrated self‑attention into a GNN for structural classification of glass‑forming liquids (yoshikawa_graph_2025), while Bapst et al., trained a GNN solely on initial particle positions to predict long‑time glass dynamics (bapst_unveiling_2020). Although these studies demonstrate the power of GNNs for property prediction and classification in glassy systems, none simultaneously address structural generation. Our approach fills this gap by learning an internal generative mechanism that bridges state‑of‑the‑art molecular‑graph generative models with physically consistent atomic reconstructions.
@@
Physical-informed Descriptors Global, physics‑informed descriptors—such as Crystallographic Information Files (CIFs) for crystals (flam-shepherdLanguageModelsCan2023; zeniMatterGenGenerativeModel2024) and SMILES for organic molecules (weininger_smiles_1988; oboyle_deepsmiles_2018; skinnider_invalid_2024)—require long‑range order and thus fail for glasses. Instead, local geometric fingerprints like smooth overlap of atomic positions (SOAP) kernels (bartok_representing_2013; de_comparing_2016; lin_expanding_2024) or atom‑centered symmetry functions (ACSF) (mudassir_systematic_2022) capture only short‑range structure and omit system‑level properties such as total potential energy. The radial distribution function (RDF) encodes short‑ and medium‑range order in disordered materials. For example, Krykunov et.al., introduced atomic property–weighted RDFs with analytic Gaussian kernels that both outperform and accelerate Bag‑of‑Bonds for atomization energy prediction (krykunov_bond_2018), and Watanabe et.al., developed an on‑the‑fly RDF‑based sampling method to flag anomalous configurations and build robust water potentials (watanabe_machine_2024). To our knowledge, we are among the first to integrate an RDF‑based loss into a variational graph model for glasses, combining RDF‑driven regularization with graph representations to capture both local geometry and global energetics.
Physical-regularized Hierarchical Graph Variational Autoencoder
@@
In this work, we introduce GlassVAE, a generative framework for disordered metallic–glass configurations that jointly captures local atomic geometry and global energetics. Our approach addresses three core tasks: (i) Reconstruction: Recover atom types, interatomic displacements and distances, and absolute positions; (ii) Energy prediction: Estimate the total potential energy of each configuration; (iii) Generation: Sample new, physically realistic atomic structures from the learned latent distribution. Formally, each molecular‐dynamics snapshot is represented by $(\mathbf{R},\mathbf{t},E)$, where $\mathbf{R}=\{\mathbf{r}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}\quad\text{(positions in a periodic box }\mathbf{L}=[L_{x},L_{y},L_{z}]\text{),}$ $\mathbf{t}$ is the one‐hot encoding of atom types, and $E$ is the total potential energy. We learn $\mathcal{E}:(\mathbf{R},\mathbf{t},E)\;\mapsto\;\boldsymbol{z},\quad\mathcal{D}:\boldsymbol{z}\;\mapsto\;(\mathbf{R},\mathbf{t}),\quad\text{and}\quad\mathcal{P}:\boldsymbol{z}\;\mapsto\;\mathbf{E}$ so that the latent variable $\boldsymbol{z}$ retains all essential physical information for faithful reconstruction, accurate energy estimation, and the generation of novel, physically plausible structures.
Graph Representation and Invariance
@@
Physical interactions in materials depend only on atomic species and their relative arrangements, not on ordering, absolute position, or orientation. Therefore, atomistic representations and predictions must be invariant under permutation, translation, and rotation—swapping identical atoms, shifting all positions by $\Delta t$, or applying any orthogonal rotation $R$ leaves interatomic distances (and thus all physical properties) unchanged. To enforce these invariance, we convert each point cloud into an attributed graph $\mathcal{G}=(V,E)$, where Nodes $i\in V$ carry one‑hot features $t_{i}$ encoding atomic species i.e., $x_{i}\in\{0,1\}$; and Edges $(i,j)\in E$ connect atoms within a cutoff distance (under periodic boundaries) and carry attributes
@@
A graph neural network with symmetric (e.g., mean) message‑passing layers aggregates these features, guaranteeing invariance to node ordering and rigid‑body transformations. We further support these invariance claims with a formal theorem.
Theorem 1 (Invariance of the Graph Representation)
Let an atomic configuration be given by positions $\mathcal{R}=\{\mathbf{r}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{3}$ and one‐hot node features $\{\mathbf{t}_{i}\}$. Construct a graph $G=(V,E)$ by connecting all pairs with $\|\mathbf{r}_{i}-\mathbf{r}_{j}\|_{2}\leq\mathrm{CUTOFF}$, and assign each edge the attribute
$\mathbf{a}_{ij}=\bigl(\Delta\boldsymbol{r_{ij}},\|\mathbf{r}_{i}-\mathbf{r}_{j}\|_{2}\bigr),\quad\text{where}\ \Delta\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}.$
Suppose the graph neural network updates and pools features using a symmetric (permutation‐invariant) aggregation function $S$ (e.g. mean or sum). Then the resulting representation is invariant to
@@
so all edge‐wise inputs to the network are identical.
@@
so all distance‐based edge features are preserved.
@@
Model Architecture
To jointly capture the intricate local geometry and global properties (e.g., energies) of disordered metallic glass configurations, we propose a hierarchical graph variational autoencoder (GlassVAE). Figure 1 sketches its architecture. Sec. Graph Representation and Invariance states how we build a graph starting from an atomic configuration (also see Sec. Algorithm, Training for more details). From graph representations, we have
Dual-path Encoder The dual-path encoder that distills two complementary summaries. The graph-level encoding applies $L$ layers of message passing to embed local neighbourhoods yielding a graph vector that two linear heads convert into the variational parameters $\mu$ and $\log\sigma^{2}$. We could then reparameterize them into a decoder head $z\sim\mu+\sigma$. In parallel, the edge path feeds all $a_{ij}$ through a residual stack of $L_{e}$ blocks and averages the refined edge features into an edge descriptor $\boldsymbol{s}$.
@@
Property Predictor A separate fusion head consumes the concatenated vector $\Phi(\boldsymbol{z}\!\parallel\!\boldsymbol{s})$ to predict the total energy $\hat{E}$, allowing the model to learn structure–property relationships directly from the latent space.
@@
$\mathcal{L}=\;\alpha_{\text{node}}\,\mathcal{L}_{\text{node}}+\alpha_{\text{edge}}\,\mathcal{L}_{\text{edge}}+\alpha_{\text{energy}}$ (1)
$\mathcal{L}_{\text{energy}}+\beta_{\text{KL}}\,D_{\text{KL}}+\alpha_{\text{RDF}}\,\mathcal{L}_{\text{RDF}}.$
@@
$\mathcal{L}_{\mathrm{node}}$ is cross‑entropy between ground‑truth one‑hot atom types and predicted class probabilities; $\mathcal{L}_{\mathrm{edge}}$ is MSE on predicted distances plus a cosine‑similarity penalty to preserve directional consistency of edge features; $\mathcal{L}_{\mathrm{energy}}$ is MSE between predicted and true potential energies; $D_{\mathrm{KL}}$ is Kullback-Leibler (KL) divergence; $\mathcal{L}_{\mathrm{RDF}}$ is $\ell_{2}$ distance between soft histograms of predicted and true pairwise distance distributions.Details about how we calculated the loss can be referred to Sec.Training.
Generation
Generation is one of the key features of GlassVAE: its structured latent space enables both random and energy‐conditioned synthesis of new, physically realistic configurations through the decoder. Here, we outline the two modes of operation. We also list the results of generation in Sec. Sampling from the Latent Space.
@@
$\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\text{and set}\ \mathbf{z}=\boldsymbol{\mu}(x)+\boldsymbol{\sigma}(x)\,\odot\,\gamma\boldsymbol{\epsilon},$
where $\odot$” denotes element-wise multiplication. This turns sampling into a differentiable operation with respect to $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$. here $\boldsymbol{\epsilon}$ is a small perturbation. Then the sampled new structures could be decoded through $\hat{x}=D_{1}(z)\ \text{and}\ \hat{a}=D_{2}(z)$.
Conditional generation. To constrain the generated energy to $E_{\min},E_{\max}]$, we could use conditional generation. Formally, we assume a standard normal prior $p(z)\sim\mathcal{N}(0,I)$. To enforce $\hat{E}(z)\in[E_{\min},E_{\max}]$, we initialize $z$ at the encoder’s posterior mean and minimize
$\mathcal{L}(z)=$ $\bigl[\max(E_{\min}-\hat{E}(z),0)\bigr]^{2}+$
$\bigl[\max(\hat{E}(z)-E_{\max},0)\bigr]^{2}+\lambda\|z\|_{2}^{2}$
@@
Theoretical Contribution of Physics-Informed Regularizer
Physics-informed loss terms (i.e. energy and RDF in our case, Eq.1) do more than improve reconstruction accuracy—they can associate the latent variables to physically admissible manifolds. Below we formalize a theoretical explanation: imposing an operator‐based regularizer $\|\mathcal{A}\boldsymbol{u}-v\|_{L^{2}}$ induces a bound on the latent test error that decays with the training error and the quadrature resolution. In other words, every unit of training progress achieved by a physics-informed term translates directly into a proportionate tightening of the latent-space generalization gap (see full proof in Sec. Theorem 2 Full Demonstration).
Functional setting. Let $\mathcal{U}=W^{s,q}(\mathbb{R}^{3})$ and $\mathcal{V}=L^{p}(\mathbb{R}^{3})$ with $1\!\leq\!p,q\!\leq\!\infty$ and $s\!\geq\!0$. Given a neural approximation $\boldsymbol{u}(\cdot;\theta)\!\in\!\mathcal{U}$, the operator $\mathcal{A}:\mathcal{U}\!\to\!\mathcal{V}$ extracts a physical observable (e.g. predicted RDF, energy density), while $v\!\in\!\mathcal{V}$ encodes the target dictated by first-principles or simulation data.
Assumption 1 - Operator stability
There exists $C_{1}>0$ such that $\|\boldsymbol{u}-\boldsymbol{u}^{\prime}\|_{\mathcal{U}}\leq C_{1}\|\mathcal{A}\boldsymbol{u}-\mathcal{A}\boldsymbol{u}^{\prime}\|_{\mathcal{V}}$ for all $\boldsymbol{u},\boldsymbol{u}^{\prime}\in\mathcal{U}$.
Assumption 2 - Quadrature accuracy